What is the equation in slop-intercept form of the line perpendicular to the segment between (1,3) and (5,5) and passing through (3,4)?
Thank You

Answer by mananth(16946)

x1 y1 x2 y2
1 3 5 5

slope m = (y2-y1)/(x2-x1)
( 5 - 3 )/( 5 - 1 )
( 2 / 4 )
m= 1/ 2
the perpendivular line will have a slope which is negative reciprocal of this segment.
so m = -2
plug m=-2 & (3,4) in y=mx+b
4=-2*3+b
5=-6+b
b=10
The equation will be y=-2x+10
m.ananth@hotmail.ca